• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Olivier

    Log entry time 00:08:51 on October 4,2002

    Entry number 87294

    keyword=Night program

    The situation tonight is that we still have a problem with the cryogenic
    fluid stability in Q1L. Ed and his team have left for the night, the hall
    is now locked, and we will receive beam during the owl shift. We will do some beam line check, harp measurement and moeller measurement. Bodo will also add to the program an energy measurement, since we have the time. But we can not use the spectrometers.

    Work in the hall will resume in the morning at 8am, and we will go to restricted access to try to fix the Q1 problem and finally start taking physics data.